# Data-Free Class-Incremental Gesture Recognition with Prototype-Guided Pseudo Feature Replay

### v1

@@ III-C Truncated Cross-Entropy for New Classes
As mentioned in [11], there exists a domain difference between real data and pseudo features generated by model inversion. When pseudo features and real data are mixed for model training, the model tends to classify the mixed samples as either real or synthetic, rather than discriminating them based on the semantic information encoded within the features. To mitigate the influence of domain differences of the classifier, we introduce an additional classifier solely dedicated to the new classes. This classifier, termed Truncated Cross-entropy for New Classes (TCE), is formulated as follows:
@@ III-D Continual Classifier Re-Training
During training, we design Continual Classifier Re-Training (CCRT) strategy, which is inspired by classifier re-training [34, 35] in long-tail learning. However, in class incremental learning, this training strategy is rarely studied. Specifically, we fix the backbone after training it on the first task and keep the backbone parameters unchanged throughout subsequent incremental tasks. For each subsequent incremental task, we only re-train the classifier to adjust the decision boundary. This strategy guarantees the stability of extracted features, preserving a consistent feature space for pseudo feature generation. Fixing the backbone allows for the acquisition of a more generalized feature representation, which aids the classifier in adapting to new tasks. It also avoids overfitting to new classes and retains feature information from old classes, helping to mitigate catastrophic forgetting. The training procedure is illustrated in the supplementary material.
@@ IV-B Experimental Setup
To assess the method, we utilize two types of test sets: one encompassing all visible categories and another consisting of current categories. For the SHREC 2017 3D dataset, after the model is trained on the initial incremental task (comprising 8 categories), we keep the backbone constant, while 1 new category is added to the visible classes during each subsequent incremental task. Similarly, for the EgoGesture 3D dataset, the backbone remains fixed once the model is trained on the first incremental class task (containing 59 categories), with 4 new categories being added during each subsequent incremental class task.
@@ IV-C Results of Gesture Recognition
Figure 4: T-SNE visualization of the deep features extracted from the SHREC 2017 3D dataset. The experiemnts of class-incremental learning are conducted with 7 tasks. (a) depicts the features of the first 8 classes from PGPFR after training on task 0, (b) shows the features of the same first 8 classes from PGPFR but after training on task 6, (c) illustrates the features of all 14 classes from PGPFR after training on task 6. In all visualizations, data points are colored according to their corresponding ground-truth labels.
@@
IV-D Ablation Studies
@@
IV-D1 Effect of PFGBP, Batch Protytypes and Sharpening
@@ IV-D3 Effect of TCE
The TCE loss, defined as $\mathcal{L}_{T}$, encourages the model to learn the semantic information of different classes rather than emphasizing the domain discrepancy between real data and pseudo features. We find that removing TCE results in inferior performance for both the mean Global Accuracy and IFM.
@@ IV-D4 Effect of CCRT
we removed the CCRT strategy for baseline, meaning that the parameters of the feature extractor was no longer fixed. The results showed a significant drop in global accuracy after removing the CCRT strategy. This indicates that without CCRT, fine-tuning the model on new tasks causes the feature extractor to overly adapt to the new class samples, gradually overshadowing the representations of the old classes, leading to overfitting on the new classes. Therefore, the CCRT strategy is crucial for providing a general and stable feature representation, effectively preventing overfitting caused by small sample sizes or different data distributions.
@@ IV-D5 Further Analysis of the Accuracies of New and Old Classes
We further analyze the roles of the components in Figure 3. Given that the accuracy for all visible classes and its disparity with the accuracy of the current classes in $\mathcal{D}^{t}$ are reported in Table III, we exhibit New Accuracy and Old Accuracy. New Accuracy denotes the accuracy of all visible classes excluding those from $\mathcal{D}^{0}$, while Old Accuracy denotes the accuracy of classes in the initial task $\mathcal{D}^{0}$. We find that, without the PFGBP and the VPR, the Old Accuracy decreases significantly during class-incremental learning, which further validates the roles of the proposed PFGBP and VPR in preventing catastrophic forgetting. Without the TCE, the Old Accuracy remains almost unchanged, as this module is solely applied to the new classes $\mathcal{D}^{t}$.
IV-D6 Visualizations
@@
We utilize t-SNE [38] to reduce the dimensionality of the learned features, allowing us to visualize them and qualitatively assess the effectiveness of our proposed method. The deep features $h=\varphi(f(x))$ are projected using t-SNE and are color-coded according to their ground-truth labels in the resulting visualization. By comparing Figure 4(a) and Figure 4(b), we observe that the first 8 classes learned from Task 0 maintain clear and distinct boundaries in the test set for Task 0. Impressively, these boundaries remain well-defined even after six incremental tasks, highlighting the strong effectiveness of our method in mitigating catastrophic forgetting. Figure 4(c) displays the features of all 14 classes after learned on Task 6, using PGPFR. Recall SHREC 2017 3D is divided into 7 tasks, the visualization reveals a clear distinction between different classes.
@@
We compare our method with the BOAT-MI [8] for hand gesture recognition of certain samples, as shown in Figure 5. Our method outperforms BOAT-MI in recognizing certain challenging samples. For instance, for the sample labeled Shake in Figure 5, our model accurately identifies it with a high confidence of 1.0, while BOAT-MI mistakenly classifies it as Swipe Left with a confidence of 0.7. Similarly, for the sample labeled Swipe Up, despite relatively lower confidence, our method correctly classifies it, whereas BOAT-MI wrongly predicts it as Expand. These results demonstrate the effectiveness of our method on these challenging gesture samples.
@@ V Conclusion
In this work, we study class-incremental gesture recognition and introduce a Prototype-Guided Pseudo Feature Replay (PGPFR). This approach primarily consists of Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay (VPR), Truncated Cross-Entropy (TCE) and Continual Classifier Re-Training (CCRT). Extensive experiments on popular gesture recognition benchmarks demonstrate the effectiveness of the proposed PGPFR as well as the components. Specifically, PFGBP and VPR collaboratively tackle the issue of catastrophic forgetting of old knowledge when learning new classes. TCE mitigates the domain difference between real data and pseudo features and improves performance for new classes. We also discover that classifier re-training enhances the performance of class-incremental learning by potentially preventing overfitting to instances of new classes. We believe that it will considerably contribute to research in both data-free class-incremental learning and gesture recognition.

### v2

@@ III-C Truncated Cross-Entropy for New Classes
As mentioned in [30], there exists a domain difference between real data and pseudo features generated by model inversion. When pseudo features and real data are mixed for model training, the model tends to classify the mixed samples as either real or synthetic, rather than discriminating them based on the semantic information encoded within the features. To mitigate the influence of domain differences of the classifier, we introduce an additional classifier solely dedicated to the new classes. This classifier, termed Truncated Cross-entropy for New Classes, is formulated as follows:
@@ III-D Continual Classifier Re-Training
During training, we design Continual Classifier Re-Training strategy, which is inspired by classifier re-training [14] in long-tail learning. However, in class incremental learning, this training strategy is rarely studied. Specifically, we fix the backbone after training it on the first task and keep the backbone parameters unchanged throughout subsequent incremental tasks. For each subsequent incremental task, we only re-train the classifier to adjust the decision boundary. This strategy guarantees the stability of extracted features, preserving a consistent feature space for pseudo feature generation. Fixing the backbone allows for the acquisition of a more generalized feature representation, which aids the classifier in adapting to new tasks. It also avoids overfitting to new classes and retains feature information from old classes, helping to mitigate catastrophic forgetting. We choose classifier re-training because it neither introduces additional parameters nor requires extra training data.
@@ IV-B Experimental Setup
To assess the method, we utilize two types of test sets: one encompassing all visible categories and another consisting of current categories. For the SHREC 2017 3D dataset, we follow the 14-gesture setting. After the model is trained on the initial incremental task (comprising 8 categories), we keep the backbone constant, while 1 new category is added to the visible classes during each subsequent incremental task. Similarly, for the EgoGesture 3D dataset, the backbone remains fixed once the model is trained on the first incremental class task (containing 59 categories), with 4 new categories being added during each subsequent incremental class task.
@@ IV-C Results of Gesture Recognition
Figure 4: Influence of the batch size in the Prototype-Guided Pseudo Feature Replay. We report the mean global accuracies and IFM of all tasks on the SHREC 2017 3D, shown in subfigures (a) and (b), respectively.
(a) Task 0 features
(b) Task 6 features (first 8 classes)
(c) Task 6 features (all 14 classes)
Figure 5: T-SNE visualization of the deep features extracted from the SHREC 2017 3D dataset. The experiments of class-incremental learning are conducted with 7 tasks. Data points are colored according to their corresponding ground-truth labels.
@@
Figure 7: Comparison of confusion matrices before and after class-incremental gesture recognition: (a) Confusion matrix after training the Task 0. (b) Confusion matrix after training all tasks. The horizontal axis represents the predicted labels, while the vertical axis represents the ground-truth labels. The color intensity reflects the predicted probability.
TABLE IV: Comparison of training computational cost and inference speed. The evaluation platform is an Intel(R) Xeon(R) Gold 5218R CPU with an NVIDIA RTX 4090 GPU. The inversion time is calculated for the Task 2 while training on the SHREC 2017.
Method Training Time (s) Inference Speed (FPS)
Total Inversion CPU GPU
BOAT-MI [1] 415260 26363 15 14
PGPFR (ours) 1680 103 15 14
IV-D Ablation Studies and Analyses
@@
IV-D1 Effect of PFGBP, Batch Class Prototypes and Sharpening
@@ IV-D3 Effect of TCE
The TCE loss, defined as $\mathcal{L}_{T}$, encourages the model to learn the semantic information of different classes rather than emphasizing the domain discrepancy between real data and pseudo features. Removing TCE results in inferior performance for both the mean Global Accuracy and IFM.
@@ IV-D4 Effect of CCRT
we removed the CCRT strategy for baseline, meaning that the parameters of the feature extractor was no longer fixed. The results showed a significant drop in global accuracy after removing the CCRT strategy. This indicates that without CCRT, fine-tuning the model on new tasks causes the feature extractor to overly adapt to the new class samples, gradually overshadowing the representations of the old classes, leading to overfitting on the new classes. The CCRT is crucial for providing a general and stable feature representation, effectively preventing overfitting caused by small sample sizes or different data distributions.
@@ IV-D5 Further Analysis of the Accuracies of New and Old Classes
We further analyze the roles of the components in Figure 3. Given that the accuracy for all visible classes and its disparity with the accuracy of the current classes in $\mathcal{D}^{t}$ are reported in Table III, we exhibit New Accuracy and Old Accuracy. New Accuracy denotes the accuracy of all visible classes excluding those from $\mathcal{D}^{0}$, while Old Accuracy denotes the accuracy of classes in the initial task $\mathcal{D}^{0}$. We find that, without the PFGBP and the VPR, the Old Accuracy decreases significantly during class-incremental learning, which further validates the roles of the proposed PFGBP and VPR in preventing catastrophic forgetting. Without the TCE, the Old Accuracy remains almost unchanged, as this module is applied to new classes $\mathcal{D}^{t}$.
IV-D6 Effect of Batch Size in Batch Class Prototypes
In Sec. III-A, we leverage batch class prototypes during online feature generation, and the default batch size is 32. Figure 4 shows the effect of batch size on performance when using batch class prototypes. We observe that the average performance first increases and then decreases as the batch size grows from 8 to 256. Although batch class prototypes improve the diversity of pseudo features for each class, a batch size that is too small prevents the prototype from adequately representing the class distribution, while an excessively large batch reduces diversity and consequently degrades performance.
IV-D7 Visualizations
@@
We utilize t-SNE to reduce the dimensionality of the learned features, allowing us to visualize them and qualitatively assess the effectiveness of our proposed method. The deep features $h=\varphi(f(x))$ are projected using t-SNE and are color-coded according to their ground-truth labels in the resulting visualization. By comparing Figure 5(a) and Figure 5(b), we observe that the first 8 classes learned from Task 0 maintain clear and distinct boundaries in the test set for Task 0. Impressively, these boundaries remain well-defined even after six incremental tasks, highlighting the strong effectiveness of our method in mitigating catastrophic forgetting. Figure 5(c) displays the features of all 14 classes after learned on Task 6, using PGPFR. Recall SHREC 2017 3D is divided into 7 tasks, the visualization reveals a clear distinction between different classes.
@@
We compare our method with the BOAT-MI [1] for hand gesture recognition of certain samples in Figure 6. Our method outperforms BOAT-MI in recognizing certain challenging samples. For instance, for the sample labeled Shake, our model accurately identifies it with a high confidence of 1.0, while BOAT-MI mistakenly classifies it as Swipe Left with a confidence of 0.7. Similarly, for the sample labeled Swipe Up, despite relatively lower confidence, our method correctly classifies it, whereas BOAT-MI wrongly predicts it as Expand. These results demonstrate the effectiveness of our method on these challenging gesture samples.
IV-D8 Failure Case Analyses
To analyze failure cases of our approach for gesture recognition. we compare the confusion matrices before and after class-incremental gesture recognition Figure 7. We find that the eight initial gesture categories can be correctly recognized with accuracies above 84%. After class-incremental learning with six new categories, three of the initial categories maintain high accuracy, while the remaining five experience drops of more than 30%. The most significant degradation occurs for swipe left, whose accuracy decreases from 89% to 34%, as most of its samples are misclassified as swipe V, a newly introduced category in Task 5.
Four of the new gesture categories achieve high accuracy, while only two new categories, namely swipe X and swipe V, show relatively low accuracies of 57% and 67%, respectively. This reduction occurs because swipe X is confused with swipe right at 14%, and swipe V is confused with swipe up at 16%. These results demonstrate the effectiveness of our approach for both new and old categories in class-incremental learning.
IV-D9 Training and Inference Efficiency Analysis
We report the training time and inference speed of our approach in Table IV. Our method only introduces additional computation during training, while the inference network remains identical to that of BOAT-MI [1]. The inference speed reaches approximately 15 FPS on both CPU and GPU platforms. However, our training time is only about 0.4% of that required by BOAT-MI.
IV-D10 Limitations
One limitation is that our pseudo feature generation relies on simple linear operations. Although this design improves efficiency, it cannot accurately approximate the true distributions of old categories. Consequently, when a new category is highly similar to an existing one, our model suffers catastrophic forgetting on that old category.
@@ V Conclusion
In this work, we study class-incremental gesture recognition and introduce a Prototype-Guided Pseudo Feature Replay (PGPFR). This approach primarily consists of Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay, Truncated Cross-Entropy and Continual Classifier Re-Training. Extensive experiments on popular gesture recognition benchmarks demonstrate the effectiveness of the proposed PGPFR as well as the components. Specifically, PFGBP and Variational Prototype Replay collaboratively tackle the issue of catastrophic forgetting of old knowledge when learning new classes. Truncated Cross-Entropy mitigates the domain difference between real data and pseudo features and improves performance for new classes. We also discover that classifier re-training enhances the performance of class-incremental learning by potentially preventing overfitting to instances of new classes. We believe that it will considerably contribute to research in both data-free class-incremental learning and gesture recognition.